# Visual Position Prompt for MLLM based Visual Grounding

### v2

@@ Abstract
Although Multimodal Large Language Models (MLLMs) excel at various image-related tasks, they encounter challenges in precisely aligning coordinates with spatial information within images, particularly in position-aware tasks such as visual grounding. This limitation arises from two key factors. First, MLLMs lack explicit spatial references, making it difficult to associate textual descriptions with precise image locations. Second, their feature extraction processes prioritize global context over fine-grained spatial details, leading to weak localization capability. To address this issue, we introduce VPP-LLaVA, an MLLM equipped with Visual Position Prompt (VPP) to improve its grounding capability. VPP-LLaVA integrates two complementary mechanisms. The global VPP overlays learnable, axis-like embeddings onto the input image to provide structured spatial cues. The local VPP focuses on fine-grained localization by incorporating position-aware queries, which suggests probable object locations. We also introduce a VPP-SFT dataset with 0.6M samples, consolidating high-quality visual grounding data into a compact format for efficient model training. Training on this dataset with VPP enhances the model’s performance, achieving state-of-the-art results on standard grounding benchmarks despite using fewer training samples compared to other MLLMs like MiniGPT-v2, which rely on much larger datasets ($\sim$21M samples). The code and VPP-SFT dataset will be available at https://github.com/WayneTomas/VPP-LLaVA upon acceptance.
@@ I Introduction
Figure 1: A visual grounding case study of MLLMs: (a) LLaVA-v1.5 outputs an inaccurate bounding box based on the given query expression. (b) When provided with a position reference, VPP-LLaVA produces a suitable result. For brevity, some text instructions are omitted.
Figure 2: An illustration of VPP-LLaVA, an MLLM-based visual grounding framework with Visual Position Prompt (VPP). We utilize the global VPP to provide a global position reference for MLLMs with foundational spatial cues. Additionally, a local VPP, serving as a local position reference, is introduced to further enhance and incorporate object spatial information. For brevity, some text instructions are omitted.
Multi-modal Large Language Models (MLLMs) [1, 2, 3, 4, 5] achieve impressive results across various image-related tasks [6, 7, 8, 9], earning considerable attention from the research community. Among these tasks, visual grounding—specifically Referring Expression Comprehension (REC)—stands out as a critical challenge [10, 11, 12, 7, 13]. Unlike pure detection tasks [14, 15, 16], visual grounding involves precisely identifying locations within an image based on free-form language expressions. It is fundamental for cognitive interactions between humans and machines, with applications such as image segmentation [17], remote sensing [18] and human-robot interaction [19].
While research [20, 21] indicates that MLLMs possess a reasonable ability for spatial understanding, there is a growing consensus that these models still require further enhancement, particularly for tasks involving precise spatial reasoning like visual grounding. As illustrated in Fig. 1 (a), the grounding results from LLaVA-v1.5, for instance, reveal notable inaccuracies (the red box indicates the predicted bounding box, while the green box represents the ground truth). Although the predicted bounding box partially covers the target object, i.e., the brown toy, it suffers from both size and shape inaccuracies, failing to align well with the object’s true boundaries. These limitations highlight the need for more effective strategies to improve spatial alignment and object localization in MLLMs. To address these challenges, some studies are investigating the integration of advanced region-level enhancement modules and larger, more comprehensive visual grounding datasets into MLLMs [13, 22, 23, 6, 24]. Other approaches are exploring the incorporation of task-specific expertise, such as converting special tokens directly into bounding boxes with additional decoder, to improve localization accuracy [25, 26].
@@
Based on these observations, we propose integrating positional references as prompts within MLLMs to improve their visual grounding capabilities. While the coordinate axis shown in Fig. 1 (b) provides a global spatial guide, we also explore the potential of leveraging object position embeddings, derived from detection models, which offer local spatial cues by capturing object locations and semantic context. These two types of Visual Position Prompt (VPP)—global and local—are complementary, with the global guide helping to establish overall spatial structure, and the local cues refining object-specific localization.
Specifically, the global VPP is initialized in an axis-like form and overlaid onto the input image, providing a global spatial reference. This enables MLLMs to more effectively align coordinate information with spatial details across the image. To capture object-specific spatial and semantic information, we introduce a local VPP, which identifies potential objects within the image. This local reference helps the decoder integrate object-level details with other features. By combining both global and local prompts, our model improves spatial alignment and boosts performance in visual grounding tasks. Notably, we achieve state-of-the-art results with a relatively small training set of about 0.6M samples from our new VPP-SFT dataset, which consolidates high-quality visual grounding data into a compact form. Compared to other MLLMs like MiniGPT-v2, which require around $\sim$21M grounding samples, our approach demonstrates superior efficiency while maintaining strong performance.
@@
  • Extensive experiments demonstrate that our method not only achieves state-of-the-art performance but also does so with a substantially smaller training dataset compared to other MLLMs, highlighting its efficiency.
@@ II-B Visual Prompt
Visual prompts are used by researchers to guide models in specific tasks [19, 47, 48], as they are more direct than text prompts [49, 50]. In the Vision-and-Language Pre-training (VLP) models, VPT [51], MaPLe [50], and CMPA [52] add several learnable tokens before the CLIP visual embeddings, transferring CLIP to few-shot classification tasks. PEVL [53] and CPT [54] adapt position-sensitive vision-language tasks like visual grounding to mask token prediction. Recently, visual prompts are introduced in large vision models like SAM [55] and its variants [56, 57] to guide image segmentation. In the field of MLLMs, many works consider the visual prompt approach for specific downstream tasks. [28] and [27] present position-guided visual prompts for GPT-4V, which highly rely on the model’s own Chain-of-Thought ability. Ferret [7] and ViP-LLaVA [58] introduce hand-drawn free-form visual prompts, such as scribbles and arrows, for building user-friendly MLLMs. [59] proposes a transferable visual prompt method that can be applied to different models on downstream tasks after training on a single model. [60] suggest incorporating external knowledge, such as segmentation masks as visual prompts for MLLMs, to enhance their visual understanding performance.
@@ III-B Global Visual Position Prompt
Inspired by [27, 59, 28], we introduce a learnable global Visual Position Prompt (VPP) $\delta_{g}\in\mathbb{R}^{3\times H\times W}$ to provide a global position reference for the MLLM, initializing it in an axis-like form.
$\delta_{g}^{i}\leftarrow X_{axis}\in\mathbb{R}^{3\times H\times W},$ (1)
where $\delta_{g}^{i}$ is the initialized global VPP and $X_{axis}$ denotes the axes image with coordinate axes along its edges, using a unit scale of 0.1. As shown in Fig. 2, this design aligns with the coordinate data format in LLaVA’s training dataset, providing the MLLM with a global reference. In contrast, we avoided adding coordinate axes outside the image, as this would misalign the image coordinates with the annotated data, leading to confusion and poor performance.
@@
$X^{gp}=\alpha\cdot\mathcal{T}_{v}(X)+(1-\alpha)\cdot\mathcal{T}_{ipt}(\delta_{g}^{i}\odot M_{w}),$ (2)
where $\alpha$ is a trade-off parameter that controls the strength of the overlaid global VPP. $\mathcal{T}v(\cdot)$ represents the visual encoder’s preprocessing transformations, such as resize and padding in CLIP. $\mathcal{T}_{ipt}$ denotes an interpolation operation that scales the global prompt to fit the size of the processed input. $M_{w}$ is a binary mask with a width of $w$ around the edges, which we use to control the visible range of the global VPP. Since we initialize the global VPP as an axis-like image with coordinates along its edges, this masking is essential, as not all pixels in it need to be visible.
@@ III-C Local Visual Position Prompt
where $F_{lp}$ denotes the generated local Visual Position Prompt, and $DETR(\cdot)$ denotes the DETR model, which includes a ResNet-101 backbone, a 6-layer transformer encoder, and a 6-layer transformer decoder. Following LLaVA, we also use a 2-layer MLP to map the features of the local prompt into the LLM’s space.
$F_{lp}^{\prime}=MLP(F_{lp}),$ (6)
@@ III-D Instruction Tuning and Data Construction
TABLE I: Training data examples in the VPP-SFT dataset. We provide two samples: the first is a training data sample for visual grounding, and the second is a region captioning sample used to maintain the language capabilities of the MLLM during downstream SFT.
[…]
@@
$F^{{}^{\prime}}=concatenate([F_{gp}^{\prime},F_{lp}^{\prime}]),$ (8)
@@
As shown in Table 3, our constructed dataset primarily derives from four main sources: a subset of LLaVA-665K [3], CB-GRD [62], CB-REF [62], and Genixer [63], containing 134,864, 264,516, 87,091, and 130,000 conversations, respectively, due to the high quality of their original data. Since our model is built on the pre-trained LLaVA model, we convert all bounding boxes and instructions to follow the format used by LLaVA. Specifically, we adapt the bounding box coordinates to the LLaVA format, pad images along their longer edges, and normalize coordinates in the $(x1,y1,x2,y2)$ form. To retain LLaVA’s original language capabilities, we incorporate a small amount of region-captioning data (e.g., CB-REF subset), formatted with instructions like: Please provide a short description for this region: xxx. We aim for the tuned MLLM not only to provide coordinates in response but also to recognize and interpret coordinates as part of the language queries. The combination of these two types of data enhances the model’s spatial awareness of object locations in images, improving its ability to perceive spatial information. A sample from our dataset is shown in Table I.
It is important to note that VPP-SFT is not just a simple fusion of other datasets. Many MLLM papers introduce fragmented datasets that are hard to reuse. VPP-SFT tries to bring these scattered visual grounding datasets together for MLLMs. Unlike other instruction-tuning datasets for grounding, which often have complex formats with lots of special position tokens, our data format is refreshingly straightforward. This simplicity makes it easier to create new datasets based on ours and allows for more direct fine-tuning of models to boost MLLM spatial awareness. As a result, our dataset is highly reusable and adaptable.
@@ IV-A Benchmarks
ReferIt. To evaluate the transferability of our model’s visual grounding capabilities to other datasets, we conduct zero-shot visual grounding experiments on the ReferIt dataset. This dataset consists of 20,000 annotated images and is notable for containing ambiguous queries (e.g., any and whole). Besides, some of these samples is labeled with inaccurate bounding boxes. Officially, it is split into three subsets for training, validation, and testing, which comprise 54,127, 5,842, and 60,103 queries, respectively.
@@ IV-B Implementation Details
Training details. We initialize our model with pre-trained LLaVA-v1.5-7B parameters and fine-tune it using the AdamW optimizer with a cosine annealing scheduler. The learning rates are set as follows: 2e-5 for the LLM (Vicuna-v1.5-7B), 2e-4 for the global VPP, 2e-5 for the local VPP generator, and 2e-4 for the projector that maps the local VPP to the LLM feature space. Unless otherwise specified, our default configuration refers to the 7B model (VPP-LLaVA-7B). Following recent works such as Cambrian-1 [21], Eagle [68], and LLaVA-NeXT [4], which demonstrate that unfreezing the visual encoder is greatly beneficial to vision-centric tasks, we adopt the same strategy, setting the learning rate of the visual encoder to 2e-6. Moreover, following Ferret, we fine-tune our model on the collected VPP-SFT dataset for 3 epochs, and the global batch size is set to 64, which takes roughly 30 hours on 8×8$\times$8 × NVIDIA A800 GPUs (80GB) for the whole training process. Additionally, the width of the binary mask $w$ and the tradeoff parameter $\alpha$ in Eq. 2 are set to 30 pixels and 0.95 by default, respectively, with more details available in the Section IV-G7.

### v4

@@ Abstract
Although Multimodal Large Language Models (MLLMs) excel at various image-related tasks, they encounter challenges in precisely aligning coordinates with spatial information within images, particularly in position-aware tasks such as visual grounding. This limitation arises from two key factors. First, MLLMs lack explicit spatial references, making it difficult to associate textual descriptions with precise image locations. Second, their feature extraction processes prioritize global context over fine-grained spatial details, leading to weak localization capability. To address these issues, we introduce VPP-LLaVA, an MLLM enhanced with Visual Position Prompt (VPP) to improve its grounding capability. VPP-LLaVA integrates two complementary mechanisms: the global VPP overlays a learnable, axis-like tensor onto the input image to provide structured spatial cues, while the local VPP incorporates position-aware queries to support fine-grained localization. To effectively train our model with spatial guidance, we further introduce VPP-SFT, a curated dataset of 0.6M high-quality visual grounding samples. Designed in a compact format, it enables efficient training and is significantly smaller than datasets used by other MLLMs (e.g.,  21M samples in MiniGPT-v2), yet still provides a strong performance boost. The resulting model, VPP-LLaVA, not only achieves state-of-the-art results on standard visual grounding benchmarks but also demonstrates strong zero-shot generalization to challenging unseen datasets. The code and dataset are available at https://github.com/WayneTomas/VPP-LLaVA.
@@ I Introduction
Figure 1: A visual grounding case study of MLLMs: (a) LLaVA-v1.5 outputs an inaccurate bounding box based on the given query expression. (b) When provided with a position reference, VPP-LLaVA produces a suitable result. The masked global VPP displays the coordinate regions using a semi-transparent gray overlay, indicating that this part is both visible and learnable. For brevity, some text instructions are omitted.
Figure 2: An illustration of VPP-LLaVA, an MLLM-based visual grounding framework with Visual Position Prompt (VPP). We utilize the global VPP to provide a global position reference for MLLMs with foundational spatial cues. Additionally, a local VPP, serving as a local position reference, is introduced to further enhance and incorporate object spatial information. In the illustrative results, the coordinate regions of the masked global VPP are highlighted with a semi-transparent gray overlay, indicating that these areas are both visible and learnable. For brevity, some text instructions are omitted.
Multi-modal Large Language Models (MLLMs) [1, 2, 3, 4, 5] achieve impressive results across various image-related tasks [6, 7, 8, 9], earning considerable attention from the research community. Among these tasks, visual grounding—specifically Referring Expression Comprehension (REC)—stands out as a critical challenge [10, 11, 12, 7, 13]. Unlike pure detection tasks [14, 15, 16], visual grounding involves precisely identifying locations within an image based on free-form language expressions. It is fundamental for cognitive interactions between humans and machines, with applications such as image segmentation [17, 18], remote sensing [19] and human-robot interaction [20].
While research [21, 22, 23] indicates that MLLMs possess a reasonable ability for spatial understanding, there is a growing consensus that these models still require further enhancement, particularly for tasks involving precise spatial reasoning like visual grounding. As illustrated in Fig. 1 (a), the grounding results from LLaVA-v1.5, for instance, reveal notable inaccuracies (the red box indicates the predicted bounding box, while the green box represents the ground truth). Although the predicted bounding box partially covers the target object, i.e., the brown toy, it suffers from both size and shape inaccuracies, failing to align well with the object’s true boundaries. These limitations highlight the need for more effective strategies to improve spatial alignment and object localization in MLLMs. To address these challenges, some studies are investigating the integration of advanced region-level enhancement modules and larger, more comprehensive visual grounding datasets into MLLMs [13, 24, 25, 6, 26]. Other approaches are exploring the incorporation of task-specific expertise, such as converting special tokens directly into bounding boxes with additional decoder, to improve localization accuracy [27, 28].
@@
Based on these observations, we propose integrating positional references as prompts within MLLMs to improve their visual grounding capabilities. While the coordinate axis shown in Fig.1 (b) provides a global spatial guide, we further incorporate object position embeddings derived from detection models to supply local spatial cues by encoding object locations and semantic context. Such integration of global and local spatial priors has also been explored in other domains [31, 32]. These two types of Visual Position Prompt (VPP)—global and local—are complementary, with the global guide helping to establish overall spatial structure, and the local cues refining object-specific localization.
Specifically, the global VPP is initialized in an axis-like form and overlaid onto the input image, providing a global spatial reference. This enables MLLMs to more effectively align coordinate information with spatial details across the image. To capture object-specific spatial and semantic information, we introduce a local VPP, which identifies potential objects within the image. This local reference helps the decoder integrate object-level details with other features. The combination of global and local VPP enables our model to better capture spatial information. Trained solely on our compact 0.6M-sample VPP-SFT dataset, which consolidates high-quality grounding annotations, the model achieves state-of-the-art performance. Furthermore, VPP-LLaVA exhibits strong zero-shot generalization on challenging unseen datasets, particularly on the newly proposed multi-granular visual grounding benchmark GSEval-BBox [33], which includes difficult samples such as part-object and multi-object grounding. Compared to other MLLMs such as MiniGPT-v2 [13] and Qwen series [34], which rely on approximately 21M grounding samples, our method achieves superior performance owing to both the VPP design and the high-quality VPP-SFT dataset.
@@
  • Extensive experiments demonstrate that VPP-LLaVA achieves state-of-the-art performance and strong zero-shot generalization with far less training data than others, demonstrating its efficiency and effective design.
@@ II-B Visual Prompt
Visual prompts are used by researchers to guide models in specific tasks [20, 53, 54, 55, 56], as they are more direct than text prompts [57, 58]. In the Vision-and-Language Pre-training (VLP) models, VPT [59], MaPLe [58], and CMPA [60] add several learnable tokens before the CLIP visual embeddings, transferring CLIP to few-shot classification tasks. PEVL [61] and CPT [62] adapt position-sensitive vision-language tasks like visual grounding to mask token prediction. Recently, visual prompts are introduced in large vision models like SAM [63] and its variants [64, 65] to guide image segmentation. In the field of MLLMs, many works consider the visual prompt approach for specific downstream tasks. [30] and [29] present position-guided visual prompts for GPT-4V, which highly rely on the model’s own Chain-of-Thought ability. Ferret [7] and ViP-LLaVA [66] introduce hand-drawn free-form visual prompts, such as scribbles and arrows, for building user-friendly MLLMs. [67] proposes a transferable visual prompt method that can be applied to different models on downstream tasks after training on a single model. [68] suggest incorporating external knowledge, such as segmentation masks as visual prompts for MLLMs, to enhance their visual understanding performance.
@@ III-B Global Visual Position Prompt
Figure 3: Axis-like images used for initializing the global VPP. (a) Default version adopted in all main experiments; (b) Internal-0.05; (c) Cross-axis-0.1; (d) External-0.1 — the latter three are variants used in the ablation study to assess the effect of different positional priors. As these are the original resolution images used in training, some text elements may appear small in print; we recommend zooming in to view finer details.
Inspired by [29, 67, 30], we introduce a learnable global Visual Position Prompt (VPP) $\delta_{g}\in\mathbb{R}^{3\times 336\times 336}$ to provide a global position reference for the MLLM. It is initialized in an axis-like form and optimized without explicit constraints, enabling flexible adaptation during training.
$\delta_{g}^{i}\leftarrow\mathcal{T}_{v}(X_{axis})\in\mathbb{R}^{3\times 336\times 336},$ (1)
where $\delta_{g}^{i}$ is the initialized global VPP and $X_{axis}$ denotes the axis-like image. $\mathcal{T}v(\cdot)$ represents the visual encoder’s preprocessing transformations, such as resize, padding and normalization in CLIP. As illustrated in Fig. 3, this design is consistent with the coordinate data format found in LLaVA’s training set, thereby providing the MLLM with a global positional reference. The axis-like RGB image used for initialization is a 3-channel image with a white background, while the axes, coordinate numbers, and tick marks are rendered in black. By default, we set a unit scale of 0.1 with axes along the edges (Fig. 3 (a)). Additional variants of the axis-like initialization image and the font size of the coordinate numbers are explored through detailed ablation studies presented in Section IV-G3.
@@
$X_{gp}=\alpha\cdot\mathcal{T}_{v}(X)+(1-\alpha)\cdot\mathcal{T}_{ipt}(\delta_{g}^{i}\odot M_{w}),$ (2)
where $X$ is the input image. $\alpha$ is a trade-off parameter that controls the strength of the overlaid global VPP. $\mathcal{T}_{ipt}$ denotes an interpolation operation that scales the global prompt to fit the size of the processed input. $M_{w}$ is a binary mask with a width of $w$ applied around the edges, which is used to control the visible region of the global VPP. Since the global VPP is initialized as an axis-like image with coordinate information along its edges, the masking mechanism is necessary because not all pixels need to be visible. As shown in Fig. 2 (global VPP module), the edges of the binary mask are set to 1, while the rest of the mask is set to 0, making the coordinates of the global VPP visible. Additionally, from the response results of VPP-LLaVA, we can see that the global VPP is marked with a semi-transparent gray overlay, indicating that this part is both visible and learnable.
@@ III-C Local Visual Position Prompt
where $F_{lp}$ denotes the generated local Visual Position Prompt, and $DETR(\cdot)$ denotes the DETR model, which includes a ResNet-101 backbone, a 6-layer transformer encoder, and a 6-layer transformer decoder. Following LLaVA-v1.5, we also use a 2-layer MLP to map the features of the local prompt into the LLM’s space.
$F_{lp}^{\prime}=MLP(F_{lp}).$ (6)
@@ III-D Instruction Tuning and Data Construction
TABLE I: Original annotation examples from four data sources. We randomly select two samples from each source for illustration. It can be seen that the formats of the four data sources are all different from each other. We unify them into the format of LLaVA-665K, and after refining and merging, we generate our VPP-SFT dataset.
Source (subset) Annotations
[…]
    {“from”: “gpt”, “value”: “The giant doughnut with white icing and red , white , and blue sprinkles.”}]
2. {[“from”: “human”, “value”: “Please provide the bounding box coordinate of the region this sentence describes: white frosting.”},
    {“from”: “gpt”, “value”: ”[0.52, 0.59, 0.82, 0.83]”}]
CB-GRD 1. [{“from”: “human”, “value”: ”Could you please tell me where is the black pants? ”},
    {“from”: “gpt”, “value”: ”I have provided the box of the black pants. <black pants:[245, 384, 283, 502]>”}]
2. {[“from”: “human”, “value”: ”Draw the red and hanging and blue jacket out. ”},
    {”from: “gpt”, “value”: “I have provided the box of the red and hanging and blue jacket. <red and hanging and blue jacket:[532, 354, 678, 570]>”}]
CB-REF 1. [{“from”: “human”, “value”: “What can you see in this <region>? <black charger:[354, 435, 451, 561]>”},
    {“from”: “gpt”, “value”: “It is a black charger.”}]
2. {[“from”: “human”, “value”: “What is <it>? <black clothing:[350, 112, 521, 153]>”},
    {“from: “gpt”, “value”: “It is a black clothing.”}]
Genixer 1.   “bbox”: [75, 13, 240, 67], “expression”: “the watch has a black strap”
2.   “bbox”: [313, 183, 431, 308], “expression”: “a road through the grass”
@@
Figure 5: Training data examples in the VPP-SFT dataset. We provide two samples: the first is a training data sample for visual grounding, and the second is a region captioning sample used to maintain the language capabilities of the MLLM during downstream SFT. The green bounding boxes marked in the images are the Ground Truth.
@@
$F^{{}^{\prime}}=concatenate([F_{gp}^{\prime},F_{lp}^{\prime}]).$ (8)
@@
As shown in Fig. 4, our constructed dataset primarily derives from four main sources: a subset of LLaVA-665K [3], CB-GRD [70], CB-REF [70], and Genixer [71], containing 134,864, 264,516, 87,091, and 130,000 conversations, respectively, due to the high quality of their original data. For each source, we provide corresponding annotation examples in Table I. As evident, the data formats vary widely. To enable consistent and efficient reuse in our training, we standardize them by converting bounding box coordinates into the LLaVA format, padding images along their longer edge, and normalizing the $(x_{1},y_{1},x_{2},y_{2})$ coordinates accordingly. Specifically, for LLaVA-665K, we extract both region captions and visual grounding conversations. In the subset of CB-GRD and CB-REF, we eliminate their special formatting schemes, preserve only referring expressions, and convert absolute coordinates (derived from original image dimensions) to normalized coordinates compatible with padded square inputs required by our model. Similarly, for Genixer, we construct conversations following grounding templates used in the LLaVA-665K and normalize the coordinate annotations to align with the LLaVA format. To retain LLaVA’s original language capabilities, we incorporate a small amount of region-captioning data (e.g., CB-REF subset), formatted with instructions like: Please provide a short description for this region: xxx. We aim for the tuned MLLM not only to provide coordinates in response but also to recognize and interpret coordinates as part of the language queries. The combination of these two types of data enhances the model’s spatial awareness of object locations in images, improving its ability to perceive spatial information. A sample from our dataset is shown in Fig. 5.
It is important to note that VPP-SFT is not just a simple fusion of other datasets. Many MLLM papers introduce fragmented datasets that are hard to reuse. VPP-SFT tries to bring these scattered visual grounding datasets together for MLLMs. Unlike other instruction-tuning datasets for grounding, which often have complex formats with lots of special position tokens, our data format is refreshingly straightforward. This simplicity makes it easier to create new datasets based on ours and allows for more direct fine-tuning of models to boost MLLM spatial awareness. Furthermore, we observe in our experiments, particularly on the zero-shot GSEval-BBox benchmark (Section IV-E), that VPP-LLaVA trained on our collected VPP-SFT dataset demonstrates impressive generalization to challenging part-object and multi-object scenarios. These results highlight the high quality and strong reusability of our dataset.
@@ IV-A Benchmarks
ReferIt. To evaluate the zero-shot of our model’s visual grounding capabilities to other datasets, we conduct zero-shot visual grounding experiments on the ReferIt dataset. This dataset consists of 20,000 annotated images and is notable for containing ambiguous queries (e.g., any and whole). Besides, some of these samples is labeled with inaccurate bounding boxes. Officially, it is split into three subsets for training, validation, and testing, which comprise 54,127, 5,842, and 60,103 queries, respectively.
GSEval-BBox. To further assess the zero-shot capabilities of our proposed method in more challenging scenarios beyond ReferIt, we evaluate it on the recently introduced GSEval-BBox [33] dataset. This dataset contains 3,715 samples and spans four fine-grained expression types: stuff, part-object, multi-object, and single-object. Among these, part-object and multi-object references are particularly challenging, requiring models to accurately localize object subregions or reason over multiple entities simultaneously. Moreover, the referring expressions in GSEval-BBox are generally longer and more complex than those in RefCOCOg, posing an additional challenge for visual grounding models.
@@ IV-B Implementation Details
Training details. We initialize our model with pre-trained LLaVA-v1.5-7B/13B parameters and fine-tune it using the AdamW optimizer with a cosine annealing scheduler. The learning rates are set as follows: 2e-5 for the LLM (Vicuna-v1.5-7B/13B), 2e-4 for the global VPP, 2e-5 for the local VPP generator, and 2e-4 for the projector that maps the local VPP to the LLM feature space. Unless otherwise specified, our default configuration refers to the 7B model (VPP-LLaVA-7B). Following recent works such as Cambrian-1 [22], Eagle [76], and LLaVA-NeXT [4], which demonstrate that unfreezing the visual encoder is greatly beneficial to vision-centric tasks, we adopt the same strategy, setting the learning rate of the visual encoder to 2e-6. Moreover, following Ferret, we fine-tune our model on the collected VPP-SFT dataset for 3 epochs, and the global batch size is set to 64, which takes roughly 30 hours on 8×8$\times$8 × NVIDIA A100 GPUs (80GB) for the whole training process. Additionally, the width of the binary mask $w$ and the tradeoff parameter $\alpha$ in Eq. 2 are set to 30 pixels and 0.95 by default, respectively, with more details available in the Section IV-G9.